What is the slope and y intercept of -2x+3y=3
Answer by geetha_rama(94)

-2x + 3y = 3
Rewriting the above in form of
y = mx + c where m is slope and c is the y intercept
3y = 2x + 3
y = (2/3)x + 1
slope = 2/3
intercept = 1